• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name C. C. Chang

    Log entry time 00:23:18 on January22,2003

    Entry number 92868

    keyword=Shift Summary

    Run summary

    Start Stop Run # Events Comments
    17:16 17:21 20467/1467/1431 22k/15k/7k
    17:27 17:32 20468/1469/1432 22k/19k/8k
    17:38 17:57 20469/1469/1437 67k/60k/34ki See events from cell wall
    17:59 18:05 20470/1470/1438 24k/21k/11k Still cell wall
    20:32 20:55 20471/1471/1439 71k/66k/39k Still no improvement

    16:30 IOCha14 hanged. Lost left arm magnet readouts. Rebooted iocha14 twice. In the process of rebooting, we Q2, D, and Q3 all tripped.
    17:15 Start first run. We also noticed that with raster off, the intrinsic spot
    size was larger than 250 um and with a halo (tail).
    17:40 With raster on, the spot size was roughly 6 x 6 instead of the nominal
    size of 4 x 4. Also notice that with empty ref cell run, there are too
    many events coming from the cell wall.
    18:00 Asked MCC to reduce raster setting from 3 to 2.5 to reduce the spot
    size to about 4 x 4 and take another run. Still see events coming cell
    wall. We decided to ask MCC to do something. Put in BeO and with
    raster off. Nilanga and J.P. went to MCC to talk to them.
    19:00 MCC is trying to tune the beam. As soon as they can get the spot
    size and halo to be resonable (?) we will take a run on empty ref cell
    with raster on to see whether there is any improvement in terms of
    wall events.
    20:00 MCC called to let us know that they could not get smaller spot size
    or halo. They will call optic expert to come and take a look.
    20:25 MCC called to give the beam back to us even though there was no
    improvement in terms of spot size or beam halo. While they are
    waiting for optic expert to come in we will take another run and see
    whether there might be any improvement. So, we ask MCC to put
    in target mask and move empty ref cell in.
    21:00 Still no improvement. MCC said the expert suggested to load other
    parameters file and will continue to try. MCC requested to put in BeO
    since they need to turn rater off.
    22:50 Surles Law, the Crew Chief (?) came to let us know that they seemed
    not to be able to reduce the spot size. They are still on the phone with
    optic guy and getting some suggestions. J. P. suggesgted that we
    should take a run with pre-buncher turned off. We checked with MCC
    and they said they already did that earlier and did not see any
    improvement. When they did that the slit remained wide. We asked
    Surles whether we could have another test run (if HJall C agrees)
    with pre-buncher off and with slit narrowed down (keeping the same
    beam intensity). He said he will let us know if such a time slot (about
    30 minutes) should become available.